A Company bought a used delivery truck for $20,000 with an estimated useful life of 4 years. The company’s accountant estimated that the truck’s residual value would be $2,000 at the end of its useful life. Compute depreciation expense for the third year of the asset’s life using the double-declining balance method.
Answer #1

Annual depreciation as per straight line=100%/4=25%/year

Hence depreciation as per double decline=2*Annual depreciation as per straight line*Beginning value of each period

Year Beginning value Depreciation Ending value
1 20,000 (2*25%*20,000)=10,000 (20,000-10000)=$10,000
2 10,000 (2*25%*10,000)=5000 (10,000-5000)=$5000
3 5000 (2*25%*5000)=$2500

Hence depreciation expense for third year=$2500